### v3

@@ 4.2 Investigating gradient conflicts between tasks and samples
Challenged paradigm: The field of MTL strongly focuses on resolving conflicts between tasks, especially from a perspective of gradient conflicts [31, 55, 20]. In computer vision, tasks are often defined on a conceptional level such as segmentation and depth (CityScapes), or recognizing multiple attributes (CelebA). However, in principle, conflicts can not only occur between tasks but also between samples within a task.
@@
We argue that in an extreme case, even recognizing a single cat in multiple images could be considered MTL. For instance, in one image, the cat could be hiding behind a plant and only revealing its eyes, requiring a neural network to recognize the cat solely based on the eyes. In other images, the cat might only reveal its paws, front of a bright window, or might be tired and curled up into a furry ball because we took so many pictures. This would require a paw, shape or fur classifier. Thus, a neural network is required to solve multiple sub-tasks to reliably recognize our cat.
@@
Approach: During the training on aforementioned datasets, we examine gradient similarity across two different setups: (1) between gradients of different tasks with respect to a single sample (inter-task), e.g., $\mathbf{g}=\nabla_{\phi}L_{0}\left(f_{\bm{\theta}}(\mathbf{x}_{i})\right)$ and $\mathbf{g^{\prime}}=\nabla_{\phi}L_{1}\left(f_{\bm{\theta}}(\mathbf{x}_{i})\right)$; and (2) between gradients corresponding to the same task but different samples within a batch (inter-sample), e.g., $\mathbf{g}=\nabla_{\phi}L_{t}\left(f_{\bm{\theta}}(\mathbf{x}_{0})\right)$ and $\mathbf{g^{\prime}}=\nabla_{\phi}L_{t}\left(f_{\bm{\theta}}(\mathbf{x}_{1})\right)$. For both setups, we compute the gradient cosine similarity and gradient magnitude similarity as well as the ratio of conflicting gradient parameters. We are aware that our comparison between samples and tasks is not direct. Nonetheless, it serves as a coarse indicator to estimate their impact during network training. Implementation details are in Appendix A3.
@@
Results: We show the evolvement of the gradient similarity measures over epochs in Fig. 4. Surprisingly, when comparing inter-sample (red line) and inter-task (blue line), we find no consistent evidence for gradient alignment conflicts (left column) to be an exclusive problem of having multiple tasks. For instance, for Cityscapes, the variation of gradient alignment is fully encapsulated within the spread we observe in inter-sample variation (task is fixed). For CelebA (LABEL:fig:gradsim_celeba), the converse seems to be mostly the case. Furthermore, the choice of network architecture and distribution of task-specific and shared parameters (SegNet vs. DeepLabV3) can have a large influence on the spread of the cosine-similarity. Both architectures have roughly a similar number of shared-parameters. However, DeepLabV3 has a higher number of task-specific parameters which seems to reduce the variance in conflicts for both inter-sample and inter-task (row one vs. two). In line with these observations, we found a similar number of conflicting gradient parameters (third column) for both inter-sample and inter-task comparisons among all experiments.
@@
For gradient magnitude similarities (middle column), we observe a clearer pattern. The similarity in magnitudes are continuously (in the mean) less pronounced for the inter-task setup compared to inter-samples (blue line is below red one in all settings). Interestingly, the relative difference between the two setups remains similar over training which justifies the choice of fixed scalar task weightings as done in [49]. Further measures can be found in Figures A8, 4.3, A9, A10 and A11.
@@
Conclusion: We find that the difficulty of MTL (inter-task and inter-sample) as opposed to STL (inter-sample only) is predominantly due to differences in gradient magnitudes. Although the problem of conflicting gradients has been typically associated with task-specific conflicts [55, 31, 20], we found that gradient alignment conflicts can actually be even more pronounced between samples. One the one hand, these observations are along the same lines as findings by Royer et al. [42] who reason that ’correcting conflicting gradients [between tasks] at every training iteration can be superfluous’. One the other hand, gradient-alignment methods in MTL could be considered not only in the context of task-specific conflicts but also for conflicts between samples.
4.3 Robustness of multi-task representations on corrupted data
In the last part of our analysis, we investigate whether features learned for multiple tasks generalize better to corrupted data compared to those learned for single tasks only.
Motivation: In his seminal paper, Caruana gives preliminary evidence that MTL provides stronger features and avoids spurious correlations (referred to better attribute selection[4]. More recently, spurious correlations have often been directly connected with robustness [13, 18]. Results from current literature on the robustness of MTL features are mixed. While MTL is stated to increase the adversarial- and noise-robustness over STL[24, 37, 53], others argue features selected by MTL could be more likely to be non-causal and, therefore, less robust [17, 2]. Here, we further challenge the paradigm whether MTL features lead to better robustness. We would like to nuance that we do not consider the transferability of representations, e.g., to new tasks, but solely focus on the claim that the MTL trained features are more robust w.r.t. different inputs.
Figure 5: Transfer to corrupted data for MTL and STL on CityScapes with DeepLabV3 and EW. We show the difference over relative performance decrease over all corruption modes averaged over five levels of severity and three runs. We color blocks if either STL or MTL can handle the respective corruption better for all metrics of a task. MTL is more robust compared to STL models for the depth task. In contrast, STL deals slightly better on the semantic segmentation task.
Table 3: Out-Of Distribution transfer on corrupted CityScapes [9] for different MTO methods on DeepLabV3. We report difference between relative performance decrease for STL and MTL averaged over all modes of corruption and all levels of severity (Eq. 6) as well as over runs for three seeds. A value lower than zero indicates a better generalization capability of the MTL backbone and vice versa.
Sem.Seg. Depth Mean
DeepLabV3 EW 0.0087 0.0017 -0.0108 -0.1826 -0.0457
[…]
RLW 0.0263 0.0084 -0.0809 -0.3595 -0.1015
IMTL 0.0172 0.0041 0.1747 0.0857 0.0704
PCGrad 0.0090 0.0019 0.0394 -0.2240 -0.0434
CAGrad 0.0224 0.0067 0.0898 -0.2841 -0.0413
Results: Figure 5 shows $\delta_{t}$ for different corruption types for a DeepLabV3 with EW on CityScapes averaged over five corruption levels and three seeds. On the semantic segmentation tasks, the STL models show a slightly lower decrease in performance on the corrupted data than MTL ($\delta_{t}>0$ more often; shaded in red), indicating that the features learned for these respective tasks can better generalize to corrupted data. In contrast, the MTL model shows significant better relative performance on the depth task ($\delta_{t}<0$ more often; shaded green). Comparing these observations to other dataset+network combinations (Sec. 4.3), we find a strong robustness of MTL features for both setups of CityScape+SegNet and NYUv2+DeepLabV3 over all tasks. However, evaluation on NYUv2+SegNet resulted on average in better performance on STL features, especially for the depth task. Furthermore, we see little evidence for general higher robustness against certain types of corruption (e.g., higher robustness against weather conditions) for either MTL or STL across all setups.
The results of other MTO methods (Secs. 4.3, 4.3 and 4.3) indicate that it depends less on the used method but more on the choice of dataset and network architecture whether some tasks would benefit from MTL for learning more robust features. Averaged absolute scores can be found in Secs. 4.3 and 4.3.
@@ 5 Conclusion and outlook
We show that common optimization methods from single task learning (STL) like the Adam optimizer are effective in MTL problems. We explain this with Adam’s partial loss-scale invariance. Next, we compare gradient conflicts during training between tasks and samples. While gradient magnitudes are a specific problem between tasks (MTL) and thus justify the need for multi-task specific methods for automatic loss weighting, we find the variability in gradient alignment to be similar between samples and tasks. Thus, we encourage a more unified viewpoint in which specific MTO methods are also considered in single-task problems and vice versa. For the ongoing debate regarding the robustness of features from MTL approaches, we find empirical evidence that multi-task features can be more robust than single-task features depending on tasks and setup. Here, future work could try to disentangle the different current perspectives from causal features [17], adversarial/noise robustness [24, 37, 53] and our results on common corruption robustness.
Beyond our work, we encourage research to improve the understanding of challenges and paradigms in MTL. For instance, our understanding of task (and sample) specific capacity allocation within a network and how best to tune it to custom requirements, is still not thoroughly understood. Often task-weights are increased to assign more importance to a task which is in contrast to tuning the learning rate per task where a smaller learning rate can be beneficial. Thus, we require further investigations and disentanglement of these two concepts.
@@ Acknowledgments
$\begin{split}\mathcal{L}&=\sum_{t}\frac{\alpha_{t}\mathcal{L}_{t}}{\alpha_{t}sg[\mathcal{L}_{t}]}=\sum_{t}\frac{\mathcal{L}_{t}}{sg[\mathcal{L}_{t}]}\end{split}$ (12)
@@
  • EW using SGD with scalings $10\cdot L_{seg}$,
  • EW using Adam EW using Adam with scalings $10\cdot L_{seg}$ and $0.01\cdot L_{dep}$ UW-O) using SGD UW-O) using SGD with scalings $10\cdot L_{seg}$ and $0.01\cdot L_{dep}$ EW using separate Adam optimizers per task EW using separate Adam optimizers per task with scalings $10\cdot L_{seg},0.01\cdot L_{dep}$ To better control for different factors of influence, we first perform the first 6 of the listed experiments with a fixed backbone, i.e., we do not update the parameters in the backbone but only in the heads. Afterward, we show all 8 experiments trained with a network where all parameters (including the backbone) are updated. This allows us to verify if our theoretical derivations regarding the (partial) loss-scaling invariance of Adam and UW-O also hold in practice, and compare this to the SGD optimizer. Note that we only care about the invariance and did not tune any hyperparameters for performance.
@@
    Corrupted variants
    For Cityscapes and NYUv2 we apply the common corruptions proposed by [16]. We use their original code [^3] https://github.com/hendrycks/robustness.
@@
    All presented results are based on performing early stopping w.r.t. $\Delta_{m}$- metric on the validation set. For this, we further trained single-task learning (STL) models for each experiment combination (dataset and network) using the respective network architecture except for the missing head(s). We trained the models using Adam and any learning rate from $\{0.01,0.005,...,0.00005\}$. The training was stopped early based on the validaton loss. Reported scores in LABEL:tab:res_nyu_segnet, LABEL:tab:res_nyu_deeplab, LABEL:tab:res_cs_segnet and LABEL:tab:res_cs_deeplab are computed as the mean of the models’ performance that were initialized with the three different seeds.
@@
    We present additional evaluation results for our comparison between optimizers for MTL. In Figure A3, we compare the $\Delta_{m}$-metric performance between the usage of Adam and SGD+mom. Sec. 4.3 shows additional parallel coordinate plots for NYUv2 and both choices of networks. In Table A2, we count for each used MTO method the number of experiment runs that are located on the Pareto front w.r.t. each setup. Best performing quantitative results for all MTOs can be found in LABEL:tab:res_cs_segnet, LABEL:tab:res_cs_deeplab, LABEL:tab:res_nyu_segnet and LABEL:tab:res_nyu_deeplab.
    We further show extended results on the toy task by Liu et al. [31] for more learning rates in LABEL:fig:cagrad_toytask_extended and LABEL:tab:cagrad_toytask_niter_full.
@@
    Figure A3: Mean $\Delta_{m}$-metric for experiments run on CityScapes and NYUv2 with SegNet and DeepLabV3. We compare the performance of the best hyperparameter setting for every MTO method using either Adam (left) or SGD+Momentum (right) (lower is better). Every MTO is associated with a different line color/style. On Cityscapes, there is a large difference for the $\Delta_{m}$ score for Adam compared to SGD+Momentum, especially for UW, IMTL, and CAGrad. Therefore, for this setup, the result depends more on the optimizer than on the MTO method. On the NYUv2 dataset this observation weakens. Adam still achieves the lowest $\Delta_{m}$ scores across different MTO methods (except for SegNet with UW and IMTL), though, besides chosing Adam, it is also important to select the appropriate MTO method.
@@
    Figure A4: Parallel coordinate plot over all experiments on NYUv2. We distinguish between experiments using SGD+mom and Adam optimizer. Experiments that reached Pareto front performance are drawn with higher saturation. Similiar to results on CityScapes in the main paper (Fig. 2) we observe a dominance of Adam albeit, here, we also have some experiments using SGD+Mom. on the overall Pareto front.

### v5

@@ 4.2 Investigating gradient conflicts between tasks and samples
Examined paradigm: The field of MTL strongly focuses on resolving conflicts between tasks, especially from a perspective of gradient conflicts [31, 56, 20]. In computer vision, tasks are often defined on a conceptional level such as segmentation and depth (Cityscapes), or recognizing multiple attributes (CelebA). However, in principle, conflicts can not only occur between tasks but also between samples within a task.
@@
We argue that in an extreme case, even recognizing a single cat in multiple images could be considered MTL. For instance, in one image, the cat could be hiding behind a plant and only revealing its eyes, requiring a neural network to recognize the cat solely based on the eyes. In other images, the cat might only reveal its paws, front of a bright window, or might be tired and curled up into a furry ball because we took so many pictures. This would require a paw, shape or fur classifier. Thus, a neural network is required to recognize multiple attributes to reliably recognize our cat. We note that this is conceptually similar to the commonly [44, 56] considered MTL dataset CelebA [35] which requires attribute detection such as wavy hair, mustache or hat, but within one image.
@@
Approach: During the training on aforementioned datasets, we examine gradient similarity across two different setups: (1) between gradients of different tasks with respect to a single sample (inter-task), e.g., $\mathbf{g}=\nabla_{\phi}L_{0}\left(f_{\bm{\theta}}(\mathbf{x}_{i})\right)$ and $\mathbf{g^{\prime}}=\nabla_{\phi}L_{1}\left(f_{\bm{\theta}}(\mathbf{x}_{i})\right)$; and (2) between gradients corresponding to the same task but different samples within a batch (inter-sample), e.g., $\mathbf{g}=\nabla_{\phi}L_{t}\left(f_{\bm{\theta}}(\mathbf{x}_{0})\right)$ and $\mathbf{g^{\prime}}=\nabla_{\phi}L_{t}\left(f_{\bm{\theta}}(\mathbf{x}_{1})\right)$. For both setups, we compute the gradient cosine similarity and gradient magnitude similarity as well as the ratio of conflicting gradient parameters. We are aware that our comparison between samples and tasks is not direct. Nonetheless, it serves as a coarse indicator to estimate their impact during network training. Implementation details are in App. A3.
@@
Results: We show the evolvement of the gradient similarity measures over epochs in Fig. 4. Surprisingly, when comparing inter-sample (red line) and inter-task (blue line), we find no consistent evidence for gradient alignment conflicts (left column) to be an exclusive problem of having multiple tasks. For instance, for Cityscapes, the variation of gradient alignment is fully encapsulated within the spread we observe in inter-sample variation (task is fixed). For CelebA, the converse seems to be mostly the case. Furthermore, the choice of network architecture and distribution of task-specific and shared parameters (SegNet vs. DeepLabV3) can have a large influence on the spread of the cosine-similarity. Both architectures have roughly a similar number of shared-parameters. However, DeepLabV3 has a higher number of task-specific parameters which seems to reduce the variance in conflicts for both inter-sample and inter-task (row one vs. two). In line with these observations, we found a similar number of conflicting gradient parameters (third column) for both inter-sample and inter-task comparisons among all experiments.
@@
For gradient magnitude similarities (middle column), we observe a clearer pattern. The similarity in magnitudes are continuously (in the mean) less pronounced for the inter-task setup compared to inter-samples (blue line is below red one in all settings). Interestingly, the relative difference between the two setups remains similar over training which justifies the choice of fixed scalar task weightings as done in [50]. Further measures can be found in Figures A6, A4 and A7.
@@
Conclusion: We find that the difficulty of MTL (inter-task and inter-sample) as opposed to STL (inter-sample only) is predominantly due to differences in gradient magnitudes. Balancing different magnitudes is tackled in the literature, e.g., [21, 50] The problem of conflicting gradients has been typically associated with task-specific conflicts [56, 31, 20], here, we find that gradient alignment conflicts can actually be even more pronounced between samples. On the one hand, these observations are along the same lines as findings by Royer et al. [42] who reason that ’correcting conflicting gradients [between tasks] at every training iteration can be superfluous’. On the other hand, gradient-alignment methods in MTL could be considered not only in the context of task-specific conflicts but also for conflicts between samples. Interestingly, previous work has explored the potential benefit of not only learning weights per task but also per sample in the dataset [49]. While our experiments show a relatively high similarity in gradient magnitude across samples and, thus, don’t motivate a sample-wise loss weighting, this could, however, be due to only little disruptive noise within data samples which has been the main motivation of [49].
@@ 5 Conclusion and outlook
We show that common optimization methods from single task learning (STL) like the Adam optimizer are effective in MTL problems. Next, we compare gradient conflicts during training between tasks and samples. While gradient magnitudes are a specific problem between tasks (MTL) and thus justify the need for multi-task specific methods for automatic loss weighting, we find the variability in gradient alignment to be similar between samples and tasks. Thus, we encourage a more unified viewpoint in which specific MTO methods are also considered in single-task problems and vice versa.
Beyond our work, we encourage to improve the understanding of challenges and paradigms specific to MTL. For instance, our understanding of task (and sample) specific capacity allocation within a network and how best to tune it to custom requirements, is still not thoroughly understood. Often task-weights are increased to assign more importance to a task which is in contrast to tuning the learning rate per task where a smaller learning rate can be beneficial. Thus, we require further investigations and disentanglement of these two concepts.
@@ Acknowledgments
$\begin{split}\mathcal{L}&=\sum_{t}\frac{\alpha_{t}\mathcal{L}_{t}}{\alpha_{t}sg[\mathcal{L}_{t}]}\\ &=\sum_{t}\frac{\mathcal{L}_{t}}{sg[\mathcal{L}_{t}]}\end{split}$ (11)
@@
  • EW using SGD with scalings $10\cdot L_{seg}$, $0.01\cdot L_{dep}$
  • EW using Adam
  • EW using Adam with scalings $10\cdot L_{seg}$ and $0.01\cdot L_{dep}$
  • UW-O) using SGD
  • UW-O) using SGD with scalings $10\cdot L_{seg}$ and $0.01\cdot L_{dep}$
  • EW using separate Adam optimizers per task
  • EW using separate Adam optimizers per task with scalings $10\cdot L_{seg},0.01\cdot L_{dep}$
To better control for different factors of influence, we first perform the first 6 of the listed experiments with a fixed backbone, i.e., we do not update the parameters in the backbone but only in the heads. Afterward, we show all 8 experiments trained with a network where all parameters (including the backbone) are updated. This allows us to verify if our theoretical derivations regarding the (partial) loss-scaling invariance of Adam and UW-O also hold in practice, and compare this to the SGD optimizer.
Note that we only care about the invariance and did not tune any hyperparameters for performance.
@@
All presented results are based on performing early stopping w.r.t. $\Delta_{m}$ metric on the validation set. For this, we further trained single-task learning (STL) models for each experiment combination (dataset and network) using the respective network architecture except for the missing head(s). We trained the models using Adam and any learning rate from $\{0.01,0.005,...,0.00005\}$. The training was stopped early based on the validation loss. Reported scores in Tabs. A6, A7, A4 and A4 are computed as the mean of the models’ performance that were initialized with the three different seeds.
@@
A3.3 Loss functions
Cityscapes
For the task of semantic segmentation, we employ a pixel-wise cross-entropy loss:
$\mathcal{L}_{CE}=\sum_{c=1}^{C}y^{c}\cdot\log(p^{c})$ (22)
where $C$ is the number of classes, $y^{c}\in\{0,1\}$ indicates the ground truth class, and $p^{c}$ is the predicted probability for class $c$ which results from computing the softmax for output logits $z^{c}$, $p^{c}=\frac{\exp(z^{c})}{\sum_{c=1}^{C}\exp(z^{c})}$. This loss is averaged over the image.
For depth estimation, we utilize the $L_{1}$ loss:
$\mathcal{L}_{depth}=\|\bm{y}-\hat{\bm{y}}\|_{1}$ (23)
where $\bm{y},\hat{\bm{y}}$ indicate ground truth and prediction, respectively. Pixels with invalid depth value in the ground truth data are ignored. It is noteworthy that these two types of losses are not balanced when used directly without modification.
NYUv2
The tasks of semantic segmentation and depth estimation are trained using the same loss functions as described for Cityscapes. In addition, we compute the cosine loss on the (normalized) surface normal maps:
$\mathcal{L}_{normal}=1-\cos{\theta}=1-\frac{\bm{y}\cdot\hat{\bm{y}}}{\|\bm{y}\|\|\hat{\bm{y}}\|}$ (24)
where $\bm{y},\hat{\bm{y}}$ are the ground truth and predicted normal maps. Similar to the Cityscapes setup, the combination of these loss functions is not balanced per default.
CelebA
To learn to predict multiple attributes, we use a binary cross entropy loss for the individual classes:
$\mathcal{L}_{CE,bin}=-[y\log(p)+(1-y)\log(1-p)]$ (25)
Although all these losses have a similar scale, their impact varies based on the difficulty of the individual tasks and the number of available samples displaying the respective attribute.
A3.4 Evaluation criteria
In this study, we primarily focus on Pareto optimal solution to acknowledge that different configurations may lead to varying preferences for the learned tasks. However, it is important to note that not every point on the Pareto front is relevant in practice, especially when one metric significantly dominates while others are close to chance level. Moreover, specific real-world applications can have a stronger, pre-defined prioritization of one or a few sub-tasks which requires a relative weighting of the tasks’ performances.
Additionally, we further employ the $\Delta_{m}$ metric which offers a simple option to directly compare the performance of two models using a single scalar. This metric further indicates the relative performance compared to the single-tasks models.
Note that for our initial toy task experiment (Sec. 4.1), we consider the original setting from [31] which optimzes the global minimum of the two loss functions.
@@
We present additional evaluation results for our comparison between optimizers for MTL. In Figure A3, we compare the $\Delta_{m}$ metric performance between the usage of Adam and SGD+mom. Fig. A4 shows additional parallel coordinate plots for NYUv2 and both choices of networks. In Table A2, we count for each used MTO method the number of experiment runs that are located on the Pareto front w.r.t. each setup. Best performing quantitative results for all MTOs can be found in App. A4, A4, A6 and A7.
We further show extended results on the toy task by Liu et al. [31] for more learning rates in Tab. A3.
@@
Figure A3: Mean $\Delta_{m}$ metric for experiments run on Cityscapes and NYUv2 with SegNet and DeepLabV3. We compare the performance of the best hyperparameter setting for every MTO method using either Adam (left) or SGD+Momentum (right) (lower is better). Every MTO is associated with a different line color/style. On Cityscapes, there is a large difference for the $\Delta_{m}$ score for Adam compared to SGD+Momentum, especially for UW, IMTL, and CAGrad. Therefore, for this setup, the result depends more on the optimizer than on the MTO method. On the NYUv2 dataset this observation weakens. Adam still achieves the lowest $\Delta_{m}$ scores across different MTO methods (except for SegNet with UW and IMTL), though, besides chosing Adam, it is also important to select the appropriate MTO method.
@@
Figure A4: Parallel coordinate plot over all experiments on NYUv2 We distinguish between experiments using SGD+mom and Adam optimizer. Experiments that reached Pareto front performance are drawn with higher saturation.
Similiar to results on Cityscapes in the main paper (Fig. 2) we observe a dominance of Adam albeit, here, we also have some experiments using SGD+Mom. on the overall Pareto front.
@@
Table A3: Number of iterations after which all seeds in toy task experiment from CAGrad [31] have reached the global minimum for different learning rates and optimizer. We show results for additional learning rates compared to the main paper. The maximum iteration number over all three seeds for each MTO method / learning rate / optimizer combination is reported. If not all seeds converged to the global minimum within 100k iteration steps, we denote it as ’-’. In several setups, EW+Adam converges fastest to the global minimum. Especially for small learning rates, CAGrad performs advantageous compared to EW. As reported in previous work, we found that PCGrad often would converge only to some point on the Pareto Front. The best and second best run for each learning rate over all MTO methods are indicated via font type.
learning rate
method 10.0 5.0 1.0 0.5 0.1 0.05 0.01 0.005 0.001*
GD EW - 103 - - - - - - -
PCGrad - - - - - - - - -
CAGrad 644 - 213 621 8,069 5,732 20,418 34,405 -
Adam EW 26 37 22 58 709 2,135 9,015 16,005 -
PCGrad 25 4,960 56 15,741 34,175 41,438 - - -
CAGrad 27 30 32 106 802 7,109 11,239 14,323 57,700
*LR used for results in [31] with Adam